A 25.0 mL sample of HCL was titrated to the endpoint with 15.0 mL of 2.0 N NaOH.What was the normality of the HCL?What was its molarity
1
Expert's answer
2013-03-13T10:05:02-0400
The chemical equation is
NaOH + HCl = NaCl + H2O

According to the equation the amount of NaOH = the amount of HCl
That's why we can write the equation

C(NaOH) x V(NaOH) = C(HCl) x V(HCl)
C(HCl) = C(NaOH) x V(NaOH)/V(HCl) = 2.0 x 15.0 / 25.0 = 1.2 N

For the acids that have only one hydrogen atom the molarity = normality

Answer:
The normality of HCl = 1.2 N
The molarity of HCl = 1.02 M